OpenStudy (helpblahblahblah):

Plz help will give medal and fan :) Which expression is equivalent to 2^4 ⋅ 2^−7? A. 1 over 2 to the power 11 B. 1 over 2 to the power 3 C. 2^3 D. 2^11

OpenStudy (nightwielder):

Multiplication means addition of the exponents. Hence: (2^4)*(2^-7)=2^(4-7)=2^-3 Therefore: 2^4*2^7=2^-3

OpenStudy (helpblahblahblah):

thxxxxxxxxx :)